How many grams are present in 5.6 litres of CO2(carbon dioxide) at STP?
1
Expert's answer
2019-06-19T14:10:28-0400

Solution: At STP 1 mol of any gas has a volume of 22.4 liters, so in 5,6 liters of CO2 there are 5.6/22.4=0.25 moles. 1 mol of CO2 has a mass of (12+16*2)=44 g, so in 0.25 moles there are 44*0.25=11 g.


Answer: 11 grams.